1. You have 75.0 cm3 of a 0.150 mol dm-3 solution of zinc sulphate (ZnSO4). What mass of zinc sulphate crystals will be left behind on evaporation of the water?
1
Expert's answer
2020-11-04T14:22:27-0500

The concentration of zinc sulphate (ZnSO4)

= 0.150 mol dm-3


Means,

1 dm3 solution contain 0.150 mol ZnSO4



Volume of the solution is = 75.0 cm3


As, 1 cm = 10–1 dm

1 cm3 = (10–1 dm)3

= 10–3 dm3


So,

Volume of the solution is = 75.0 cm3

= 75.0 × 10–3 dm3


So,

1 dm3 solution contain 0.150 mol ZnSO4

75.0 × 10–3 dm3 solution contain

= 75.0 × 10–3 × 0.150 mole

= 11.25×10–3 mole

So, the given amount of solution contain 11.25×10–3 mole ZnSO4.

On evaporation 11.25×10–3 mole of ZnSO4 will left behind.




Molar mass of ZnSO4 = 161.47 g/mole


So,

1 mole ZnSO4 = 161.47 g/mole ZnSO4

11.25×10–3 mole ZnSO4

= (161.47 × 11.25×10–3) g

= 1.817 g ZnSO4


Hence,

1.817 g zinc sulphate (ZnSO4) crystals will be left behind on evaporation of the water